If a,b,c be the sum of n terms of three AP's whose first terms are unity and common difference are in H.P. then n=
a.(2ac+ab+bc)/(a+c-2b)
b. (2ac-ab-bc)/(a+c-2b)
c. (2ac-ab-bc)/(a+c+2b)
d. (2ac-ab+bc)/(a+c+2b)

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear Vaibhav

let common difference are d1,d2 ,d3

a = n/2[2 +(n-1)d1]  or    1/d1 = n(n-1)/2(a-n)

similerly     1/d2 = n(n-1)/2(b-n)

                1/d3 = n(n-1)/2(c-n)

d1 ,d2 ,d3 are in HP

2/d2 = 1/d1 +1/d2

 n(n-1)/(b-n) = n(n-1)/2(a-n)   +  n(n-1)/2(c-n)

   1/(b-n) = 1/2(a-n)   +  1/2(c-n)

n= (2ac -ab-bc)/(a+c-2b)


Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you
the answer and detailed solution very quickly.

 We are all IITians and here to help you in your IIT JEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin